OpenStudy (anonymous):

Help please! Can someone help me find x? (7x+.1)+(10x-.2)=5

OpenStudy (anonymous):

x=0.3

OpenStudy (anonymous):

No, I need help working it out, not just the answer!

OpenStudy (leilanilane):

The Golden Rule of Math: What you do to one side you must do to the other. PEMDAS (7x + 1) + (10x - 2) = 5 - 1 - 1 (7x) + (10x - 2) = 4 + 2 + 2 (7x) + (10x) = 6 7x + 10x = 6 Can you do the rest?

OpenStudy (anonymous):

Absolutely, thanks so much!

OpenStudy (anonymous):

dfrtgr edededededededededededededededededededededededededededededededededededededededededededed